OpenStudy (anonymous):

11 1/3 divided by 2 5/6=

OpenStudy (anonymous):

4

OpenStudy (anonymous):

\[\frac{11\frac{1}{3}}{2\frac{5}{6}}\]?

OpenStudy (anonymous):

11 1/3 is same as 34/3 and 2 5/6 is the same as 17/6 so 34/3 divided by 17/6 is just 34/3 times 6/17 and then you just multiply across and simplify to get your answer.